OpenStudy (anonymous):

If f(x) = x2 + 1, find f(3). Everytime I do this problem I get 8, but I know that's wrong because my answer choices are: 10 7 4 2

OpenStudy (abb0t):

If you plug in 3 for x, you get (3)(3) = 9. add one to that, and what do you get?

OpenStudy (abb0t):

\[(3)^2 = (3)(3) \] similarly \[(3)^7 = (3)(3)(3)(3)(3)(3)(3) \] and so on.

OpenStudy (anonymous):

ohhhh... okay. :) Now I get it. Instead of doing (3)(3), I was doing 3^2... duh!!! lol... THANKS!!

OpenStudy (anonymous):

and then I wasn't adding the 1 on... lol.